# Generalized Recurrence Criteria for Classes of Open Quantum Walks

### v1

@@ 3.1 Faithfulness of Density Operators
1. Let $\rho\in\mathcal{S}(\mathfrak{h}_{i})$ be faithful. By the compactness of $\mathcal{S}(\mathfrak{h}_{i}),$ there exists $c>0$ such that $\rho>c\tau,$ thus
$\begin{split}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=&\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\\ >&c\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)(I\otimes\lvert i\rangle\langle i\rvert)\right]\\ =&c\sum_{n=0}^{\infty}p_{ii;\tau}(n)\\ =&\infty.\end{split}$
2. Let $\rho\in\mathcal{S}(\mathfrak{h}_{i})$ be non-faithful. By the Spectral Theorem, $\rho$ can be written as
$\rho=\sum_{x=1}^{n}\lambda_{x}\lvert x\rangle\langle x\rvert,$ (2)
where the vectors $\lvert x\rangle$ are the eigenvectors of $\rho$ with corresponding eigenvalues $\lambda_{x}$. Since $\rho$ is non-faithful, there is at least one null eigenvalue and the remainder of eigenvalues are non-negative summing $1$ Thus, (2) can be rewritten as
$\rho=\sum_{x\in S}\lambda_{x}\lvert x\rangle\langle x\rvert,\quad S\varsubsetneq\{1,\ldots,n\}.$
Take a sequence of positive numbers $(\alpha_{r})_{r\in R},$ where $R:=\{1,\ldots,n\}/S\neq\emptyset,$ whose sum is $1$
Defining
$\rho_{X}=\sum_{x\in S}\frac{\lambda_{x}}{2}\lvert x\rangle\langle x\rvert+\sum_{x\in R}\frac{\alpha_{x}}{2}\lvert x\rangle\langle x\rvert=\sum_{x=1}^{n}\frac{\tilde{\alpha}_{x}}{2}\lvert x\rangle\langle x\rvert,\quad\tilde{\alpha}_{x}=\begin{cases}\lambda_{x},&\mbox{if }x\in S\\ \alpha_{x},&\mbox{if }x\in R\end{cases},$
we get by the first item that $\lvert i\rangle$ is $\rho_{X}$-recurrent, since $\rho_{X}$ is faithful.
Now, define
$\rho_{Y}=\sum_{x\in R}\alpha_{x}\lvert x\rangle\langle x\rvert,$
which is a non-faithful density operator. Since $2\rho_{X}=\rho+\rho_{Y},$ we obtain by linearity the equality
$\sum_{n=0}^{\infty}p_{ii;\rho}(n)+\sum_{n=0}^{\infty}p_{ii;\rho_{Y}}(n)=\sum_{n=0}^{\infty}\left(p_{ii;\rho}(n)+p_{ii;\rho_{Y}}(n)\right)=2\sum_{n=0}^{\infty}p_{ii;\rho_{X}}(n).$
The sum on the right hand diverges once $\lvert i\rangle$ is $\rho_{X}$-recurrent, implying that at least one of the sums on the left-hand side diverges. Therefore, $\lvert i\rangle$ is $\rho$-recurrent or $\rho_{Y}$-recurrent, both being non-faithful.
3. Let $\tilde{\rho}$ be one non-faithful density given by item (2), then we have
$\tilde{\rho}=\sum_{j}\lambda_{j}\lvert e_{j}\rangle\langle e_{j}\rvert,$
where each $\lambda_{j}\geq 0$ is an eigenvalue of $\tilde{\rho}$ with corresponding eigenvector $\lvert e_{j}\rangle.$ If $\tilde{\rho}$ has only one eigenvector, then we are done. If $\tilde{\rho}$ has at least two eigenvectors, then
$\infty=\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]=\sum_{j,\lambda_{j}\neq 0}\left(\lambda_{j}\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\lvert e_{j}\rangle\langle e_{j}\rvert\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\right),$
where at least one of the members of the sum must be infinity, and some $\lambda_{j}$ is positive. This completes the proof.
@@ 3. Let $\tilde{\rho}$ be one non-faithful density given by item (2), then we have
Let $\Phi$ be an OQW on some set of vertices $V$ and $i\in V,$ and $\dim(\mathfrak{h}_{i})<\infty.$ If $\lvert i\rangle$ is $\tau$-transient for some faithful $\tau\in\mathcal{S}(\mathfrak{h}_{i}),$ then $\lvert i\rangle$ is transient.
@@
Let $\Phi$ be a homogeneous OQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{S}(\mathfrak{h}),$ then
@@
    $\Phi$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{S}(\mathfrak{h})$;
@@
    there exists a non-faithful density operator $\tilde{\rho}$ such that $\Phi$ is $\tilde{\rho}$-recurrent;
  3. 3.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that $\Phi$ is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unitary vector of $\mathbb{C}^{d}.$
@@ 3.2 Enclosures and Invariant States
Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every positive $\sigma\in L^{1}(\mathfrak{h}),$ we have
@@
Following this property and using the homogeneity of OQWs induced by a coin, we split $\mathcal{S}(\mathfrak{h})$ into two subsets $\mathcal{S}(\mathcal{Y})$ and $\mathcal{S}(\mathcal{X})$ in such a way that the walk is transient for $\mathcal{S}(\mathcal{Y})$ and recurrent for $\mathcal{S}(\mathcal{X}),$ where $\mathcal{Y}$ is a enclosure of $\mathcal{L}:$
@@
Consider an OQW induced by a coin $(L,B,R)$ of dimension $d<\infty$. We are in one (and only one) of the following situations:
@@
  • $(L,B,R)$ is recurrent with respect to all densities except for a convex set $\mathcal{S}(\mathcal{Y}).$ In this case, $\mathcal{Y}$ is an enclosure of $\mathcal{L},$
    $\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X}$
@@
Let us suppose that $(L,B,R)$ is neither recurrent nor transient. By Proposition 3.1, there exist unitary vectors $\lvert u\rangle,\lvert v\rangle\in\mathfrak{h}$ such that the walk is $\lvert u\rangle\langle u\rvert$-recurrent and $\lvert v\rangle\langle v\rvert$-transient. Let $\mathcal{Y}$ be the subset of $\mathfrak{h}$ such that $\mathcal{S}(\mathcal{Y})$ is the set of densities of $\mathcal{B}(\mathfrak{h})$ in which $(L,B,R)$ is transient, thus we pick $\rho\in\mathcal{S}(\mathcal{Y})$ and we can affirm that $\rho_{n}\in\mathcal{S}(\mathcal{Y}),$ where $(X_{n},\rho_{n})_{n}$ are the quantum trajectories of the OQW. Indeed, if we had $\rho_{n}\in\mathcal{S}(\mathfrak{h})\setminus\mathcal{S}(\mathcal{Y})$ for some $n,$ then we could refresh the walk and the vertex $\lvert X_{n}\rangle$ would be $\rho_{n}$-recurrent, contradicting the hypothesis of homogeneity, thus we must have that $\mathcal{Y}$ is an enclosure of $\mathcal{L}.$ Therefore, take $\mathcal{X}=\mathcal{Y}^{\perp}$ and the trichotomy holds.
@@
Consider a Lazy OQW induced by a coin $(L,B,R)$ of dimension $d<\infty.$ If the auxiliary map $\mathcal{L}$ admits only one invariant state $\rho_{\infty},$ then
$\frac{X_{n}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}m,\qquad\text{where }\qquad m=\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)-\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right),$
@@
This result can be obtained by using a Doob decomposition as described in [8, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
@@
Theorem 3.6 (Chung-Fuchs Theorem for OQWs with ergodic auxiliary map).
Consider an OQW induced by a coin $(L,B,R)$ such that $\mathcal{Y}$ is an enclosure Suppose that the initial density operator is $\rho\in\mathcal{S}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{S}(\mathcal{Y}).$
@@
The proof is analogous to [9, Theorem 7] and we will proceed by remarking on how it can be generalized. Firstly, the maximum appearing in [9, Lemma 6] can be taken in $\mathcal{D}=\mathcal{S}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{S}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{S}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [9, Lemma 6] can be reduced to $\mathcal{S}(\mathcal{Y}).$ Therefore, for the same reason, the density operator considered in [9, Theorem 7] belongs to $\mathcal{S}(\mathcal{Y}).$
@@
Let us consider an OQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic. Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
@@
We suppose that the walk is neither recurrent nor transient. In this case, the walk is recurrent for some density and transient for all densities in some $\mathcal{S}(\mathcal{Y}),$ where $\mathcal{Y}$ is an enclosure (Proposition 3.4). Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{S}(\mathcal{Y}).$
If $\;\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.5, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{S}(\mathcal{Y})$ by Theorem 3.6, which is a contradiction.
@@
3.3 General Recurrence Criteria for OQWs induced by a coin $(L,B,R)$ of dimension 2
The following Proposition is a generalisation of [9, Proposition 12, item (3)]. We can prove it in an analogous demonstration given by [9], however, we would like to do it in terms of the non-ergodicity of the auxiliary map induced by this coin.
Proposition 3.8.
Consider an OQW induced by a coin $(L,B,R)$ of dimension 2 such that $L,B,R$ share two unitary linearly independent eigenvectors. Then for any $\lvert u\rangle$ orthogonal to one of these two eigenvectors, we have that $\lvert u\rangle$ is also a common eigenvector of $L,B,R.$
Proof.
Suppose $L,B,R$ share two unitary and linearly independent unitary eigenvectors $\lvert v\rangle,\lvert w\rangle,$ then $\tau_{1}=\lvert v\rangle\langle v\rvert$ and $\tau_{2}=\lvert w\rangle\langle w\rvert$ are distinct invariant states of the auxiliary map $\mathcal{L},$ thus $\mathcal{L}$ is not ergodic. This means that there are two minimal enclosures $\mathcal{Y}_{1},\mathcal{Y}_{2},$ and thus $\mathcal{Y}_{1}\perp\mathcal{Y}_{2}.$ The dimension of the coin gives that $\mathcal{Y}_{1}$ and $\mathcal{Y}_{2}$ are one-dimensional and therefore there exist $\lvert u_{1}\rangle,\lvert u_{2}\rangle\in\mathbb{C}^{2}$ such that $\mathcal{Y}_{1}=\mbox{span}(\lvert u_{1}\rangle)$ and $\mathcal{Y}_{2}=\mbox{span}(\lvert u_{2}\rangle).$ Thereby $\lvert u_{1}\rangle$ and $\lvert u_{2}\rangle$ are common orthonormal eigenvectors of $L,B,R.$
$\square$
@@ 3.3 General Recurrence Criteria for OQWs induced by a coin $(L,B,R)$ of dimension 2
Consider an OQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
@@
(2) If $L,B,R$ have two linearly independent eigenvectors in common, $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ unitaries, then they satisfy $\lvert u_{1}\rangle\bot\lvert u_{2}\rangle,$ we can put
@@
$(2.3)\quad|l_{i}|=|r_{i}|\mbox{ and }|l_{j}|\neq|r_{j}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert\mbox{ and it is}$
recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma_{j}.$
@@
Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Proposition 3.8 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert,$ where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent. Since it is valid for any common eigenvector, the walk is recurrent by Proposition 3.1.
Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert,\;\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert$ where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or $>$) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$. Since the walk is transient for two distinct eigenvectors, the walk is transient by Proposition 3.1.

### v2

@@ 3.1 Faithfulness of Density Operators
Let $\rho\in\mathcal{D}(\mathfrak{h})$ be faithful. Since the state space $\mathcal{D}(\mathfrak{h})$ is compact in the trace-norm topology, the function
$f(\tau)=\inf\{c>0\mid\rho>c\tau\}$
is continuous and thus attains its minimum and maximum. In particular, there exists $c>0$ such that
$\rho>c\tau,\quad\forall\rho\in\mathcal{D}(\mathfrak{h}).$
thus
$\begin{split}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=&\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\\ >&c\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)(I\otimes\lvert i\rangle\langle i\rvert)\right]\\ =&c\sum_{n=0}^{\infty}p_{ii;\tau}(n)\\ =&\infty,\end{split}$
giving the first item.
For the second one, let
$\tau=\sum_{j}\lambda_{j}\lvert e_{j}\rangle\langle e_{j}\rvert,$
where each $\lambda_{j}>0$ is an eigenvalue of $\tau$ with corresponding eigenvector $\lvert e_{j}\rangle.$ Then
$\infty=\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]=\sum_{j}\left(\lambda_{j}\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\lvert e_{j}\rangle\langle e_{j}\rvert\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\right),$
where at least one of the members of the sum must be infinity. This completes the proof.
@@
Let $\Phi$ be an OQW on some set of vertices $V,$ $i\in V,$ and $\dim(\mathfrak{h})<\infty.$ If $\lvert i\rangle$ is $\tau$-transient for some faithful $\tau\in\mathcal{D}(\mathfrak{h}),$ then $\lvert i\rangle$ is transient.
@@
Let $\Phi$ be a HOQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
@@
    $\Phi$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{D}(\mathfrak{h})$;
@@
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that $\Phi$ is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unit vector of $\mathbb{C}^{d}.$
@@ 3.2 Enclosures and Invariant States
Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every $\sigma\in\mathcal{D}\mathfrak{h}),$ we have
@@
Following this property and using the homogeneity of HOQWs induced by a coin, we split $\mathfrak{h}$ into two subsets $\mathcal{Y}$ and $\mathcal{X}$ in such a way that the walk is transient for $\mathcal{D}(\mathcal{Y})$ and recurrent everywhere else, with $\mathcal{Y}$ being an enclosure of the quantum channel with Kraus operators $L,B,R.$
@@
Consider a HOQW induced by a coin $(L,B,R)$ of dimension $d$. We are in one (and only one) of the following situations:
@@
  • $(L,B,R)$ is recurrent with respect to all densities except for a convex set $\mathcal{D}(\mathcal{Y}).$ In this case, $\mathcal{Y}$ is an enclosure of $\mathcal{L},$
    $\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X}$ (3)
@@
We claim $(L,B,R)$ is $\rho$-transient if and only if it is $\rho_{n}$-transient for every $n,$ where $\rho_{n}$ is the density at the $n$-th step of the walk.
Indeed, let us suppose the walk starts at vertex $\lvert 0\rangle$ with initial density $\rho.$ For instance, if the first jump is to $\lvert-1\rangle,$ the density, after this first jump, is
$\rho_{1}=\frac{L\rho L^{*}}{\mathrm{Tr}(L\rho L^{*})}.$
Since the walk is homogeneous, it is $\rho$-transient if and only if it is $\rho_{1}$-transient. After $n$ jumps, the density is
$\rho_{n}=\frac{J_{\pi}\rho J_{\pi}^{*}}{\mathrm{Tr}(J_{\pi}\rho J_{\pi}^{*})},$ (4)
where $J_{\pi}$ is a concatenation of $n$ operators, each one belonging to $\{L,B,R\},$ proving the claim, once this is valid for any $n=1,2,3,\ldots$
Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be null, and by definition $\mathcal{Y}$ is an enclosure.
Let us suppose that $(L,B,R)$ is neither recurrent nor transient. By Theorem 3.1, there exist unit vectors $\lvert u\rangle,\lvert v\rangle\in\mathfrak{h}$ such that the walk is $\lvert u\rangle\langle u\rvert$-recurrent and $\lvert v\rangle\langle v\rvert$-transient. Let $\mathcal{Y}$ be the subspace of $\mathfrak{h}$ such that $\mathcal{D}(\mathcal{Y})$ is the set of densities of $\mathcal{B}(\mathfrak{h})$ in which $(L,B,R)$ is transient, thus we pick $\rho\in\mathcal{D}(\mathcal{Y})$ and we can affirm that $(L,B,R)$ is $\rho_{n}$-transient for every $n.$ Finally, take $\mathcal{X}=\mathcal{Y}^{\perp}$ and the trichotomy holds.
@@
Consider a Lazy HOQW induced by a coin $(L,B,R)$ of dimension $d.$ If the auxiliary map $\mathcal{L}$ admits only one invariant state $\rho_{\infty},$ then
$\frac{X_{n}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}m,\qquad\text{where }\qquad m=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)-\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right),$
@@
This result can be obtained by using a Doob decomposition as described in [10, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
@@
Theorem 3.6 (Chung-Fuchs Theorem for HOQWs with ergodic auxiliary map).
Consider a HOQW induced by a coin $(L,B,R),\;\mathcal{L}$ its auxiliary map and $\mathcal{Y}$ an enclosure. Suppose that the initial density operator is $\rho\in\mathcal{D}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{D}(\mathcal{Y}).$
@@
The proof is analogous to [11, Theorem 7] and we will proceed by remarking on how it can be generalized. Firstly, the maximum appearing in [11, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{D}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [11, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [11, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case. Moreover, the lazy case holds as an extension of this proof, since all the results work for homogeneous OQWs on $\mathbb{Z}$.
@@
The following theorem presents the first recurrence criterion for non-irreducible HOQWs of arbitrary finite dimension. This result extends [11, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
@@
Let us consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic. Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
@@
We suppose that the walk is neither recurrent nor transient. In this case, the walk is recurrent for some density and transient for all densities in some $\mathcal{D}(\mathcal{Y}),$ where $\mathcal{Y}$ is an enclosure (Proposition 3.4). Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{D}(\mathcal{Y}).$
If $\;\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.5, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.6, which is a contradiction.
@@
3.3 General Recurrence Criteria for HOQWs induced by a coin $(L,B,R)$ of dimension 2
The following theorem is due to [4].
@@ 3.3 General Recurrence Criteria for HOQWs induced by a coin $(L,B,R)$ of dimension 2
For a finite quantum channel $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$, the following are equivalent:
i.
$\mathcal{L}$ is ergodic;
ii.
$\mathcal{L}$ does not have two invariant subspaces $S_{1}\neq\{0\}$ and $S_{2}\neq\{0\}$ such that $S_{1}\cap S_{2}=\{0\};$
iii.
$\mathcal{L}$ has a minimal invariant subspace, that is, a subspace $S\neq\{0\}$ such that $S\subset S^{\prime}$ for every invariant subspace $S^{\prime}\neq\{0\}.$
Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [11, Theorem 17] to encompass the Lazy case. This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
Theorem 3.9.
Consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
@@
(2) If $L,B,R$ have two linearly independent eigenvectors in common, $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ unit, then they satisfy $\lvert u_{1}\rangle\bot\lvert u_{2}\rangle,$ we can put
@@
$(2.3)\quad|l_{i}|=|r_{i}|\mbox{ and }|l_{j}|\neq|r_{j}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert\mbox{ and it is recurrent}$
[…]
@@
Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Theorem 3.8 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert,$ where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent. If $(L,B,R)$ was not transient, then we would be in third situation of proposition 3.4, with
$\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X},\quad\text{where}\quad\text{dim}(\mathcal{Y})\geq 1,\;\;\text{ and }\;\;\text{dim}(\mathcal{X})=2,$
a contradiction. Therefore, we are in the first situation of proposition 3.4 and the walk is recurrent.
Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert,$ and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert,$ where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or $>$) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$. Since the walk is transient for two distinct eigenvectors, the walk is transient by corollary 3.2.